Mir Status Report 14 April 1997

NASA MIR 4 mission, MD 93-96 (Edited)
Date: 14 April 1997 / Mission Day 93
Mir Station System/Subsystem Activities
KOB-2 Loop Leak Repair (entire crew)

Mission Science Payload Activities: QUELD Sample E-13;DOZA - TLD Processing

Stowage/Transfer
The crew reported the failed solid oxygen generator as well as the cut-out segments of the thermal pipes have been packed in ziplock bags in preparation for their return on STS-84.

The crew also mentioned that the 'Raketa' vacuum cleaner they have been using throughout their increment is failing. The crew is unsure if the black old vacuum cleaner they found onboard is functioning. The crew recommended manifesting a replacement ASAP and suggested STS-84. The ground gave no response to this suggestion.

Mir Systems

The Elektron system has been working continuously since Saturday evening. Its O2 production is not quite enough for three crewmembers, so the O2 is being supplemented with Progress O2 as required. This procedure is described in on-board documentation. They hope to maintain this mode of operation until Shuttle arrives with the new Elektron.

The Vosdukh system is working well. It's efficiency (with respect to removing moisture) will be improved by installing a fan to blow some cooling air onto the system.

Urine Processing System - There are two parts to this system. One for collection, and the other for pumping for transfer and preservation. The pumping portion isn't working. At first, the ground thought the problem was a bad pump. This was removed and replaced, but the problem did not go away. Now, they believe the automated pumping system is receiving a false signal from somewhere. The ground is generating a fault detection scheme for the crew, but since this isn't a major problem it won't be worked for some time. For now, urine is being pumped into an EDV, then into a tank in the Progress.

While continuing the searching for the VGK leak, the crew found a relatively large amount of water behind a panel in the Kvant 1 module. The water was located near the gyrodynes in an optical sensor shroud. Since this was later and not coolant, the ground believes that this is condensate collecting at a cool place and is not a cause for concern. The crew was instructed to clean up the area and press on with leak detection.

EVA Scheduling - Prior to Progress, a committee recommended that if the repairs could not be completed by the 15 APR, the EVA should be cancelled. This committee will meet this week to status the repairs and make a recommendation on the EVA. The schedule for the 15th was based on ground estimates for the time required to effect the repairs. These estimates did not account for leaks other than the heat exchanger (which has been patched around) or for the extensive work involved in finding the materials in Progress and moving all of the 'trash' that has been accumulating in Kvant 1 over the years.

LiOH canisters - There are four 'old' and nine 'new' useable canisters onboard. Ten 'new' were delivered by Progress. One of these was used immediately. The 'old' canisters absorb moisture at the beginning of their life, so they prefer to use these for de-humidifying the Soyuz. The 'new' canisters do not absorb water.

Fire Extinguishers - The crew reported that there are three absolutely full, six 'cocked/partially used' and one empty fire extinguishers onboard. The crew does not know how much remains in the partially used fire extinguishers. Since four new ones came up on the Progress, the ground has instructed the crew to put four of the partially used extinguishers in the Progress, and then evenly distribute the new extinguishers throughout the station. (No specific instructions were given on what to do with the empty extinguisher but it seems logical to assume that this one will also be moved to the Progress.)

Repair of the MOTsA (the failed component preventing the ability to downlink MIPS) will be scheduled for the April 15th and 16th.

The MOST team received a phone call from Houston today asking about a Mir reboost scheduled for tomorrow, April 15th. From talking with the shift flight director we learned that one IS in the tomorrow's plan. Two times are currently being logged for the reboost to occur: 13:47 - 13:57 DMT and 15:23 - 15:34 DMT.

Environmental Parameters (Information from 13 April)

Parameter                             Min.   Max.   Nominal Range 

Pressure (mmHg) Base Block 747 753 660 - 850
ppO2 (mmHg) 158 160 140 - 200
ppCO2 (mmHg) 4.4 5.1 up to 8.0
ppH20 (mmHg) 12.7 13.4 8 - 14
Humidity (percentage) 37 40 30 - 75
Temperature (0C) base block (BB) 30.7 31.5 18 - 28
Temp (0C) docking/transfer module 21.6 22.1 18 - 28
Temp (0C) work area Krystall 25.4 26.0 18 - 28
Temp (0C) sleep area Krystall 26.7 27.0 18 - 28
Temp (0C) Spektr 25.1 25.6 18 - 28
Temp (0C) Kvant-2 20.2 21.2 18 - 28



Planning

MD 94
KOB-2 Loop Leak Repair (entire crew)
Condensate Recovery System (CRS) IFM (Mir 23)
QUELD Sample E-26 (NASA 4)

MD 95
KOB-2 Loop Leak Repair (entire crew)
Condensate Recovery System (CRS) IFM (Mir 23)
Air Sampling (Mir 23)
Grab Sample Container (NASA 4)
QUELD Samples E-14 & 27 (NASA 4)
Formaldehyde Monitoring (NASA 4)

  • submit to reddit


Loading



Humans to Mars Summit 2019
Military Space Situational Awareness April 1 - April 2, 2019, London, UK
MilSatCom USA conference will return to Washington from the 26th – 27th June 2019
Space Access Conference 2019
Mission To Methonē by Les Johnson - Baen Books




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on April 14, 1997 12:39 PM.

Mir Status Report 13 April 1997 was the previous entry in this blog.

Letter From Frank Culbertson to Valery Ryumin Regarding Flight Readiness Review is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.